# On Phase Unwrapping via Digital Wavefront Sensors

### v1

@@ 1 Introduction
In this paper, we consider a new class of phase unwrapping methods, which are based on mathematical observations of the working principles of certain types of wavefront sensors (WFSs). These sensors are typically used to measure optical wavefront aberrations, and are an integral part of adaptive optics (AO) systems used, e.g., in astronomy or medical imaging. Within the commonly used framework of Fourier optics, wavefront aberrations are usually modelled as exponential factors of the form $e^{i\phi}$ in the Fourier plane of the optical system. Using various optical elements, such as lenslet arrays or pyramidal prisms, wavefront sensors obtain implicit measurements of $\phi$ (e.g. its gradient). Using these measurements the full wavefront aberration $\phi$ is then computed via suitable reconstruction algorithms. Crucially, since wavefront aberrations only enter in these sensors via the exponential factor $e^{i\phi}$, the obtained measurements are the same independently of whether $\phi$ contains $2\pi$ jumps or not. This leads directly to the following key idea for phase unwrapping: One can take a wrapped phase $\phi_{w}$ and digitally propagate it through a suitable wavefront sensor, using standard tools of geometrical optics. This provides virtual wavefront measurements, from which one computes $\phi$ using any reconstruction algorithm for the corresponding wavefront sensor. Since these algorithms are designed to produce smooth wavefront reconstructions, the result of this procedure will then be an unwrapped phase $\phi$. The fact that these algorithms are typically optimized with respect to speed and reconstruction quality for aberrations with certain properties then also directly translates to wrapped phases with similar properties.
The aim of this paper is to present the idea of phase unwrapping using digital wavefront sensors outlined above in a detailed and mathematically rigorous way. In particular, we focus on phase unwrapping using digital Shack-Hartmann and Fourier-type wavefront sensors, which are commonly employed in adaptive optics both in astronomy and ophthalmology. The latter is a very broad class of sensors, and includes for example the well-known 3- and 4-sided pyramid wavefront sensors (PWFS). Both the Shack-Hartmann and Fourier-type wavefront sensors have been successfully used for reconstructing wavefront aberrations resulting from atmospheric turbulence, and thus their digital versions should be well suited for unwrapping atmospheric turbulence profiles. This is relevant e.g. in free-space optical communication (FSOC), where such turbulence profiles need to be unwrapped repeatedly and accurately. Hence, the numerical experiments presented in this paper will have a strong focus on this relevant application problem. In particular, we will see that our approach matches, and in some cases outperforms, the reconstruction quality of currently used state-of-the-art methods, and this at a comparable computational cost.
@@ 2 Physical background on wavefront sensors
In this section, we review some necessary background on WFSs, which are a key component of current and upcoming AO systems. In general, AO systems consist of one or more WFSs and deformable mirrors (DMs), and are designed to perform real-time corrections of dynamic wavefront aberrations. The aberrations are corrected by converting wavefront measurements obtained by the WFSs into DM actuator commands such that after reflection on the DMs, an incoming wavefront is restored to its non-aberrated (typically planar) form. For the determination of optimal DM actuator commands, precise phase measurements across the entire wavefront are essential. Hence, reconstructing a wavefront from WFS data and calculating DM actuator commands are therefore two crucial inverse problem.
WFSs have been commonly applied in astronomical AO systems correcting for wavefront aberrations caused by atmospheric turbulence [39, 41]. For ground-based telescopes, AO systems are employed to perform diffraction-limited observations. With the new generation of extremely large telescopes (ELTs), the development of advanced AO systems has recently gained much attention by researchers. Additionally, AO systems are implemented to enable free-space optical communication (FSOC), such as optical laser communications between spacecraft and ground-based stations. There, the goal is to equip future space missions with a communication rate orders of magnitude greater than existing channels. Furthermore, in ophthalmic AO similar instruments are employed for the early detection of anomalies and diseases in medical diagnostics. Recently, AO systems have also been applied for high- or super-resolution imaging of biological structure, and function in a wide range of microscopic modalities [54].
@@ 2.1 The Shack-Hartmann wavefront sensor
The Shack-Hartmann wavefront sensor (SH-WFS) is one of the most commonly used WFSs in AO [6, 33, 34]. It consists of a quadratic array of small lenses, also called lenslets, through which incoming light has to pass before reaching a CCD photon detector located in their focal plane. The detector itself is subdivided into a regular grid of so-called subapertures, and each lenslet focuses a part of the incoming light onto one of the subapertures; cf. Figure 2.1. Now assume that the incoming light originates from a distant point source such as an astronomical or artificial star, or a distant laser, and that no aberrations are present. In this case, the incoming wavefront of the light source is plane, and thus each lenslets focuses the light onto a (diffraction-limited) point onto its corresponding subaperture on the CCD detector; cf. Figure 2.1(left). On the other hand, if due to aberrations the wavefront is not plane, then these points are in fact shifted away from the center, and their shape is blurred; cf. Figure 2.1(middle). The SH-WFS effectively measures the change in position of the centers of mass of of these blurred points with respect to the centers of the subapertures; cf. Figure 2.1(right). These relative shifts can mathematically be connected to the average gradient (slope) of the wavefront aberration on each subapertures, from which an approximation of the incoming wavefront can then be computed numerically [39]. This is a non-trivial inverse problem, for which many different reconstruction approaches have been proposed; see e.g. [57, 43, 42]. These typically leverage physical priors such as smoothness or statistical assumptions on the wavefront aberrations, and are usually optimized in terms of numerical efficiency to satisfy stringent real-time requirements [46, 49].
@@ 2.2 Fourier-type wavefront sensors
Figure 2.2: Schematic depiction of Fourier-type WFS measurement system; cf. [21].
Next, we consider the large class of nonlinear Fourier-type WFSs [10, 22], which are included in the instrument design of many current and upcoming AO systems. Fourier-type wavefront sensing is based on optical Fourier filtering with a suitable optical element, such as a multi-facet glass prism, located in the focal plane; cf. Figure 2.2. The optical element splits the electromagnetic field into several (potentially overlapping) beams, each of which representing a differently filtered image of the entrance pupil. The number of beams after the optical element corresponds to the number of its facets dividing the light. The intensity patterns are then measured by a camera in the conjugate pupil plane. Especially in astronomical AO systems, an additional modulation component is often added to the Fourier-type WFS setup due to nonlinearity issues. This means that the electromagnetic field is moved around the apex of the optical element in order to improve the light distribution to all faces of the optical element [44], influencing both system complexity, cost-effectiveness, and sensitivity [37, 52].
@@
The key component of a Fourier-type WFS is the optical element in its focal plane. Different realisations of Fourier-type WFSs correspond to different optical elements which split the light in the focal plane, e.g., an $n$-sided pyramidal prism. The most common example in AO instruments is the $4$-sided glass pyramidal prism, typically referred to as the pyramid WFS (PWFS) [35, 36]. The prism splits the electromagnetic field into $4$ beams that are imaged in the conjugate pupil plane. Moreover, $2$-sided (roof WFS), $3$-sided, and $6$-sided prisms, as well as the extension of the idea to an infinite number of faces (cone WFS) have been investigated in more detail for astronomical AO [4, 5, 7, 52]. A Fourier-type WFSs for which the optical element does not have the shape of a prism is the Zernike WFS. It has a small circular hole in the center of the mask of the optical element [53]. The iQuad WFS is another Fourier-type WFS. Its focal plane is divided into $4$-quadrants around the origin, i.e., it has a Cartesian structure where each quadrant is phase shifted with its two neighbours [9]. As we will discuss in Section 3.4, the different optical elements can mathematically be characterized via their optical transfer functions (OTFs) [10], examples of which are given in Table 3.1 below. Visualizations of the OTFs corresponding to the 3-sided WFS, cone WFS, iQuad WFS, and different PWFSs and roof WFSs are depicted in Figure 2.3. Note that the apex angle of the prism influences the separation distance between the intensity patterns in the conjugate pupil plane (see Figure 2.3, first three columns from the left).
@@ 3.2 Key ideas and generic unwrapping algorithm
This means that when the phase $\phi$ and its wrapped form $\phi_{w}$ are considered as wavefront aberrations in the optical system corresponding to Figure 3.1, their effect on the outgoing electric field $E^{m}$ is equivalent. Consequently, any WFS using the electric field $E^{m}$ as an input for measuring the wavefront aberration is unable to distinguish between the non-wrapped and wrapped phases $\phi$ and $\phi_{w}$, respectively. Crucially, though, most wavefront reconstructors are optimized for reconstructing aberrations with certain smoothness properties, in particular those designed for the SH and Fourier-based WFSs in astronomical applications. Hence, the result of applying such a reconstructor to $E^{m}(\boldsymbol{x},\phi_{w})$ essentially yields a smooth unwrapped phase $\phi_{r}$. Additionally, the closer the properties of $\phi$ conform to those underlying a given WFS reconstruction algorithm, the better an approximation of $\phi$ the unwrapped phase $\phi_{r}$ will be.
@@
As the name suggests, Algorithm 1 is more of a general concept than a directly implementable algorithm. In particular, the concrete realization of Steps 2 and 3 will strongly impact the overall complexity and performance of the resulting method. Two possible choices, based on both SH- and Fourier-based WFSs, are considered below.
@@ 3.3 The digital Shack-Hartmann wavefront sensor
In this section, we consider the specific choice of the SH-WFS in Step 2 of Algorithm 1. In particular, we discuss how corresponding WFS sensor measurements can be calculated completely digitally, following the mathematical framework derived in [19].
As discussed in Section 2.1, the SH-WFS works via a lenslet array, each lenslet of which focuses the incoming light of a distant point source onto a corresponding subaperture of a CCD detector. Mathematically, this can be modelled by splitting the aperture domain $\Omega$ in (3.2) into a number of smaller subdomains $\Omega_{j,k}\subset\Omega$, defined by
@@
Note that we implicitly assume that the domain $\Omega$ can be split evenly into the subdomains $\Omega_{j,k}$ for a certain, potentially infinite, set of indices $\Lambda$. Next, let $\chi_{j,k}$ denote the indicator functions of $\Omega_{j,k}$ and define
@@
As with the actual SH-WFS, the number and thus the size of the subapertures $\Omega_{j,k}$ has a direct impact on the overall reconstruction quality also in the digital case. On the one hand, if the number of subapertures is too small, and thus their relative size too large, then the reconstructed wavefront will be comparatively smoother than the original wavefront, leading to insufficient resolution of finer details. On the other hand, if too many and thus too small apertures are chosen, then the reconstruction becomes susceptible to noise. See the results presented in Figure 2 for an illustrative example. Hence, a proper balance between the number and size of the subapertures needs to be struck to address this inherent ill-posedness effect. Doing so in an algorithmic way is rather difficult, requiring estimates of both the amount of noise and the expected smoothness of the original wavefront aberration. While not entirely circumventing these issues, the use of additional regularization in the wavefront reconstruction step, or alternatively the use of an “averaged” digital SH-WFS, can be beneficial here [19].
@@ 3.4 Digital Fourier-type wavefront sensors
In this section, we consider the choice of Fourier-based WFSs in Step 2 of Algorithm 1. As we shall see below, these WFSs have a simple mathematical description in terms of Fourier optics, and thus the computation of digital WFS measurements required in this step is even more straightforward than for the digital SH-WFS considered above.
First, compare the two-lens imaging system depicted in Figure 3.1 with the schematic depiction of a Fourier-type WFS measurement in Figure 2.2. The main difference is that instead of the wavefront aberration $\phi$ and the aperture $\Omega$, in Figure 2.2 there is an optical element characterized by an OTF. However, these two settings are closely related: As in Figure 3.1 let $E$ denote the incident field entering the Fourier-type WFS, and let $E^{m}$ denote the filtered field at the conjugate plane. Then, following the Fourier-optics description of this measurement system setting (cf. [21, 10, 17]), we obtain that
@@
4-sided pyramid x-Roof WFS 3-sided pyramid WFS
@@
where $\phi$ are the wavefront aberrations the be reconstructed, and the domain $\Omega$ corresponds to the aperture of the imaging system. This is physically realized if an additional lens is places in front of the imaging system in Figure 2.2, and if the incoming light stems from a sufficiently distant point source. In this case, (3.10) becomes
@@
For the reconstruction of the aberration $\phi$ from the intensity measurements $I_{\psi}$, a number of different reconstruction algorithms have been proposed. Most of these depend heavily on the shape of the chosen optical element, i.e., on the concrete form of $\psi$, and are often based on linearizations of (3.12). An overview of reconstruction methods for the most commonly used 4-sided pyramid WFS can e.g. be found in [44]. Additionally, AI-based methods such as Deep Optics PWFS have recently been proposed [51]. For arbitrary choices of $\psi$, including those which do not correspond to a pyramidal prism but can e.g. be realized by a spatial light modulator, the Nonlinear Pyramid Extension (NOPE) [21, 20] has been proposed.
Summarizing the above considerations, we arrive at the following digital Fourier-based WFS phase unwrapping algorithm as a special case of the generic Algorithm 1.
Algorithm 3 (Digital Fourier-based WFS phase unwrapping algorithm).
@@
    Reconstruct the wavefront aberration $\phi$ from the intensity measurements $I^{m}_{\psi}$ using any wavefront reconstruction algorithm for Fourier-based WFSs such as NOPE.
@@
Among the Fourier-based WFSs, the 4-sided pyramid WFS is the most commonly used. In [45], the Preprocessed Cumulative Reconstructor with Domain Decomposition (PCuReD), based on CuReD for the SH-WFS discussed above, is suggested as a suitable and highly efficient reconstructor, which gives rise to the following algorithm.
Algorithm 4 (Digital 4-sided pyramid WFS phase unwrapping algorithm).
@@
    Compute the intensity measurements $I^{m}_{\psi}$ via (3.12) with $\phi=\phi_{w}$ and choosing $\psi$ corresponding to the 4-sided pyramid WFS in Table 3.1 for some $c>0$, i.e.,
@@
Additional extensions of the Fourier-based WFS unwrapping algorithms are possible by incorporating the physical concept of modulation. Especially in astronomical AO systems, modulation of the electric field is often added to the Fourier-type WFS setup. By moving around the apex of the optical element, the light distribution to all faces of the optical element is improved, helping with certain nonlinearity issues [44].
@@
$E^{m}_{\psi}(\boldsymbol{x},\phi,t)=\mathcal{F}^{-1}\left(e^{i\psi(\boldsymbol{\xi})}\mathcal{F}\left(\chi_{\Omega}e^{-i(\phi+\varphi^{\,\operatorname{mod}}_{t}}\right)(\boldsymbol{\xi})\right)(\boldsymbol{x})\,.$ (3.13)
@@
where $T$ denotes the duration of one full modulation period. Other modulation scenarios are also possible [44]. Hence, one can again follow the concept for phase unwrapping discussed above, i.e., computing $\tilde{I}^{m}_{\psi}(\boldsymbol{x},\phi_{w})$ and using an appropriate wavefront reconstructor to obtain an unwrapped phase $\phi_{r}$. This then leads to a complete new class of modulated Fourier-based WFS phase unwrapping algorithms.

### v2

@@ 1 Introduction
In this paper, we consider a new class of phase unwrapping methods, which are based on mathematical observations of the working principles of certain types of wavefront sensors (WFSs). These sensors are typically used to measure optical wavefront aberrations, and are an integral part of adaptive optics (AO) systems used e.g. in astronomy or medical imaging. Within the commonly used framework of Fourier optics, wavefront aberrations are usually modelled as exponential factors of the form $e^{i\phi}$ in the Fourier plane of the optical system. Using various optical elements, such as lenslet arrays or pyramidal prisms, wavefront sensors obtain implicit measurements of $\phi$ (e.g. its gradient). Using these measurements the full wavefront aberration $\phi$ is then computed via suitable reconstruction algorithms. Crucially, since wavefront aberrations only enter in these sensors via the exponential factor $e^{i\phi}$, the obtained measurements are the same independently of whether $\phi$ contains $2\pi$ jumps or not. This leads directly to the following key idea for phase unwrapping: One can take a wrapped phase $\phi_{w}$ and digitally propagate it through a suitable wavefront sensor, using standard tools of geometrical optics. This provides virtual wavefront measurements, from which one computes $\phi$ using any reconstruction algorithm for the corresponding wavefront sensor. Since these algorithms are designed to produce smooth wavefront reconstructions, the result of this procedure will then be an unwrapped phase $\phi$. The fact that these algorithms are typically optimized with respect to speed and reconstruction quality for aberrations with certain properties then also directly translates to wrapped phases with similar properties.
The aim of this paper is to present the idea of phase unwrapping using digital wavefront sensors outlined above in a detailed and mathematically rigorous way. In particular, we focus on phase unwrapping using digital Shack-Hartmann and Fourier-type wavefront sensors, which are commonly employed in adaptive optics both in astronomy and ophthalmology. The latter is a very broad class of sensors, and includes for example the well-known 3- and 4-sided pyramid wavefront sensors (PWFS). Both the Shack-Hartmann and Fourier-type wavefront sensors have been successfully used for reconstructing wavefront aberrations resulting from atmospheric turbulence, and thus their digital versions should be well suited for unwrapping atmospheric turbulence profiles. This is relevant e.g. in free-space optical communications (FSOC), where such turbulence profiles need to be unwrapped repeatedly and accurately. Hence, the numerical experiments presented in this paper will have a strong focus on this relevant application problem. In particular, we will see that our approach matches, and in some cases outperforms, the reconstruction quality of currently used state-of-the-art methods, and this at a comparable computational cost.
@@ 2 Physical background on wavefront sensors
In this section, we review some necessary background on WFSs, which are a key component of current and upcoming AO systems. In general, AO systems consist of one or more WFSs and deformable mirrors (DMs), and are designed to perform real-time corrections of dynamic wavefront aberrations. The aberrations are corrected by converting wavefront measurements obtained by the WFSs into DM actuator commands such that after reflection on the DMs, an incoming wavefront is restored to its non-aberrated (typically planar) form. For the determination of optimal DM actuator commands, precise phase measurements across the entire wavefront are essential. Hence, reconstructing a wavefront from WFS data and calculating DM actuator commands are two crucial inverse problem.
WFSs have been commonly applied in astronomical AO systems correcting for wavefront aberrations caused by atmospheric turbulence [39, 41]. For ground-based telescopes, AO systems are employed to perform diffraction-limited observations. With the new generation of extremely large telescopes (ELTs), the development of advanced AO systems has recently gained much attention by researchers. Additionally, AO systems are implemented to enable FSOC, such as optical laser communications between spacecraft and ground-based stations. There, the goal is to equip future space missions with a communication rate orders of magnitude greater than existing channels. Furthermore, in ophthalmic AO similar instruments are employed for the early detection of anomalies and diseases in medical diagnostics. Recently, AO systems have also been applied for high- or super-resolution imaging of biological structure, and function in a wide range of microscopic modalities [54].
@@ 2.1 The Shack-Hartmann wavefront sensor
The Shack-Hartmann wavefront sensor (SH-WFS) is one of the most commonly used WFSs in AO [6, 33, 34]. It consists of a quadratic array of small lenses, also called lenslets. Through these incoming light has to pass before reaching a CCD photon detector located in their focal plane. The detector itself is subdivided into a regular grid of so-called subapertures, and each lenslet focuses a part of the incoming light onto one of the subapertures; cf. Figure 2.1. Now assume that the incoming light originates from a distant point source such as an astronomical or artificial star, or a distant laser, and that no aberrations are present. In this case, the incoming wavefront of the light source is plane, and thus each lenslet focuses the light onto a (diffraction-limited) point onto its corresponding subaperture on the CCD detector; cf. Figure 2.1(left). On the other hand, if due to aberrations the wavefront is not plane, then these points are in fact shifted away from the center, and their shape is blurred; cf. Figure 2.1(middle). The SH-WFS effectively measures the change in position of the centers of mass of these blurred points with respect to the centers of the subapertures; cf. Figure 2.1(right). These relative shifts can mathematically be connected to the average gradient (slope) of the wavefront aberration on each subaperture, from which an approximation of the incoming wavefront can then be computed numerically [39]. This is a non-trivial inverse problem, for which many different reconstruction approaches have been proposed; see e.g. [57, 43, 42]. These typically leverage physical priors such as smoothness or statistical assumptions on the wavefront aberrations, and are usually optimized in terms of numerical efficiency to satisfy stringent real-time requirements [46, 49].
@@ 2.2 Fourier-type wavefront sensors
Figure 2.2: Schematic depiction of a Fourier-type WFS measurement system; cf. [21].
Next, we consider the large class of nonlinear Fourier-type WFSs [10, 22], which are included in the instrument design of many current and upcoming AO systems. Fourier-type wavefront sensing is based on optical Fourier filtering with a suitable optical element, such as a multi-facet glass prism, located in the focal plane; cf. Figure 2.2. The optical element splits the electromagnetic field into several (potentially overlapping) beams, each of which representing a differently filtered image of the entrance pupil. The number of beams after the optical element corresponds to the number of its facets dividing the light. The intensity patterns are then measured by a camera in the conjugate pupil plane. Especially in astronomical AO systems, an additional modulation component is often added to the Fourier-type WFS setup due to nonlinearity issues. This means that the electromagnetic field is moved around the apex of the optical element in order to improve the light distribution to all faces of the optical element [44], influencing system complexity, cost-effectiveness, and sensitivity [37, 52].
@@
The key component of a Fourier-type WFS is the optical element in its focal plane. Different realisations of Fourier-type WFSs correspond to different optical elements which split the light in the focal plane, e.g. an $n$-sided pyramidal prism. The most common example in AO instruments is the $4$-sided glass pyramidal prism, typically referred to as the pyramid WFS (PWFS) [35, 36]. The prism splits the electromagnetic field into $4$ beams that are imaged in the conjugate pupil plane. Moreover, $2$-sided (roof WFS), $3$-sided, and $6$-sided prisms, as well as the extension of the idea to an infinite number of faces (cone WFS) have been investigated in more detail for astronomical AO [4, 5, 7, 52]. A Fourier-type WFSs for which the optical element does not have the shape of a prism is the Zernike WFS. It has a small circular hole in the center of the mask of the optical element [53]. The iQuad WFS is another Fourier-type WFS. Its focal plane is divided into $4$-quadrants around the origin, i.e., it has a Cartesian structure where each quadrant is phase shifted with its two neighbours [9]. As we will discuss in Section 3.4, the different optical elements can mathematically be characterized via their optical transfer functions (OTFs) [10], examples of which are given in Table 3.1 below. Visualizations of the OTFs corresponding to the 3-sided WFS, cone WFS, iQuad WFS, and different PWFSs and roof WFSs are depicted in Figure 2.3. Note that the apex angle of the prism influences the separation distance between the intensity patterns in the conjugate pupil plane (see Figure 2.3, first three columns from the left).
@@ 3.2 Key ideas and generic unwrapping algorithm
This means that when the phase $\phi$ and its wrapped form $\phi_{w}$ are considered as wavefront aberrations in the optical system corresponding to Figure 3.1, their effect on the outgoing electric field $E^{m}$ is equivalent. Consequently, any WFS using the electric field $E^{m}$ as an input for measuring the wavefront aberration is unable to distinguish between the non-wrapped and wrapped phases $\phi$ and $\phi_{w}$, respectively. Crucially, though, most wavefront reconstructors are optimized for reconstructing aberrations with certain smoothness properties, in particular those designed for the SH and Fourier-type WFSs in astronomical applications. Hence, the result of applying such a reconstructor to $E^{m}(\boldsymbol{x},\phi_{w})$ essentially yields a smooth unwrapped phase $\phi_{r}$. Additionally, the closer the properties of $\phi$ conform to those underlying a given WFS reconstruction algorithm, the better an approximation of $\phi$ the unwrapped phase $\phi_{r}$ will be.
@@
As the name suggests, Algorithm 1 is more of a general concept than a directly implementable algorithm. In particular, the concrete realization of Steps 2 and 3 will strongly impact the overall complexity and performance of the resulting method. Two possible choices, based on both SH- and Fourier-type WFSs, are considered below.
@@ 3.3 The digital Shack-Hartmann wavefront sensor
In this section, we consider the specific choice of the SH-WFS in Step 2 of Algorithm 1. In particular, we discuss how corresponding WFS measurements can be calculated completely digitally, following the mathematical framework derived in [19].
As discussed in Section 2.1, the SH-WFS works via a lenslet array, each lenslet of which focuses the incoming light of a distant point source onto a corresponding subaperture of detector. Mathematically, this can be modelled by splitting the aperture domain $\Omega$ in (3.2) into a number of smaller subdomains $\Omega_{j,k}\subset\Omega$, defined by
@@
Note that we implicitly assume that the domain $\Omega$ can be split evenly into the subdomains $\Omega_{j,k}$ for a certain, potentially infinite, set of indices $\Lambda$. Next, let $\chi_{j,k}$ denote the characteristic functions of $\Omega_{j,k}$ and define
@@
As with the actual SH-WFS, the number and thus the size of the subapertures $\Omega_{j,k}$ has a direct impact on the overall reconstruction quality in the digital case. On the one hand, if the number of subapertures is too small, and thus their relative size too large, then the reconstructed wavefront will be comparatively smoother than the original wavefront, leading to insufficient resolution of finer details. On the other hand, if too many and thus too small apertures are chosen, then the reconstruction becomes susceptible to noise. See the results presented in Figure 4.2 for an illustrative example. Hence, a proper balance between the number and size of the subapertures needs to be struck to address this inherent ill-posedness effect. Doing so in an algorithmic way is rather difficult, requiring estimates of both the amount of noise and the expected smoothness of the original wavefront aberration. While not entirely circumventing these issues, the use of additional regularization in the wavefront reconstruction step, or alternatively the use of an “averaged” digital SH-WFS, can be beneficial here [19].
@@ 3.4 Digital Fourier-type wavefront sensors
In this section, we consider the choice of Fourier-type WFSs in Step 2 of Algorithm 1. As we shall see below, these WFSs have a simple mathematical description in terms of Fourier optics, and thus the computation of digital WFS measurements required in this step is even more straightforward than for the digital SH-WFS considered above.
First, compare the two-lens imaging system depicted in Figure 3.1 with the schematic depiction of a Fourier-type WFS in Figure 2.2. The main difference is that instead of the wavefront aberration $\phi$ and the aperture $\Omega$, in Figure 2.2 there is an optical element characterized by an OTF. However, these two settings are closely related: As in Figure 3.1 let $E$ denote the incident field entering the Fourier-type WFS, and let $E^{m}$ denote the filtered field at the conjugate plane. Then, following the Fourier-optics description of this measurement system setting (cf. [21, 10, 17]), we obtain that
@@
4-sided PWFS x-Roof WFS 3-sided PWFS
@@
where $\phi$ are the wavefront aberrations to be reconstructed, and the domain $\Omega$ corresponds to the aperture of the imaging system. This is physically realized if an additional lens is placed in front of the imaging system in Figure 2.2, and if the incoming light stems from a sufficiently distant point source. In this case, (3.10) becomes
@@
For the reconstruction of the aberration $\phi$ from the intensity measurements $I_{\psi}$, a number of different reconstruction algorithms have been proposed. Most of these depend heavily on the shape of the chosen optical element, i.e., on the concrete form of $\psi$, and are often based on linearizations of (3.12). An overview of reconstruction methods for the most commonly used 4-sided PWFS can e.g. be found in [44]. Additionally, AI-based methods such as Deep Optics PWFS have recently been proposed [51]. For arbitrary choices of $\psi$, including those which do not correspond to a pyramidal prism but can e.g. be realized by a spatial light modulator, the NOnlinear Pyramid Extension (NOPE) [21, 20] has been proposed.
Summarizing the above considerations the following digital Fourier-type WFS phase unwrapping algorithm is proposed as a special case of the generic Algorithm 1.
Algorithm 3 (Digital Fourier-type WFS phase unwrapping algorithm).
@@
    Reconstruct the wavefront aberration $\phi$ from the intensity measurements $I^{m}_{\psi}$ using any wavefront reconstruction algorithm for Fourier-type WFSs such as the NOPE.
@@
Among the Fourier-type WFSs, the 4-sided PWFS is the most commonly used. In [45], the Preprocessed Cumulative Reconstructor with Domain Decomposition (PCuReD), based on CuReD for the SH-WFS discussed above, is suggested as a suitable and highly efficient reconstructor, which gives rise to the following algorithm.
Algorithm 4 (Digital 4-sided PWFS phase unwrapping algorithm).
@@
    Compute the intensity measurements $I^{m}_{\psi}$ via (3.12) with $\phi=\phi_{w}$ and choosing $\psi$ corresponding to the 4-sided PWFS in Table 3.1 for some $c>0$, i.e.,
@@
Additional extensions of the Fourier-type WFS unwrapping algorithms are possible by incorporating the physical concept of modulation. Especially in astronomical AO systems, modulation of the electric field is often added to the Fourier-type WFS setup. E.g. by moving around the apex of the optical element, the light distribution to all faces of the optical element is improved, helping with certain nonlinearity issues [44].
@@
$E^{m}_{\psi}(\boldsymbol{x},\phi,t)=\mathcal{F}^{-1}\left(e^{i\psi(\boldsymbol{\xi})}\mathcal{F}\left(\chi_{\Omega}e^{-i(\phi+\varphi^{\,\operatorname{mod}}_{t})}\right)(\boldsymbol{\xi})\right)(\boldsymbol{x})\,.$ (3.13)
@@
where $T$ denotes the duration of one full modulation period. Other modulation scenarios are also possible [44]. Hence, one can again follow the concept for phase unwrapping discussed above, i.e., computing $\tilde{I}^{m}_{\psi}(\boldsymbol{x},\phi_{w})$ and using an appropriate wavefront reconstructor to obtain an unwrapped phase $\phi_{r}$. This then leads to a complete new class of modulated Fourier-type WFS phase unwrapping algorithms.